A plate having an area of 0.6 m2
is sliding down an inclined plane
at 30° to the horizontal with a velocity of 0.36 m/s. There is a
cushion of fluid 1.8mm thick between the plane and the plate. Find
the viscosity of the fluid if the weight of the plate is 280N?
1
Expert's answer
2020-08-28T09:32:54-0400

"F=\eta\frac{\Delta v}{\Delta x}A\to \eta=\frac{F}{\frac{\Delta v}{\Delta x}A}="


"=\frac{mg\sin30°}{\frac{\Delta v}{\Delta x}A}=\frac{280\sin30°}{\frac{0.36}{1.8\cdot10^{-3}}\cdot0.6}=1.17(Pa\cdot s)"